# A Review of LLM-Assisted Ideation

### v2

@@ 6.2 Qualitative Analysis
inductive [xu2024jamplate_G3], [suh2024luminate_G13], [kang2024biospark_G14], [blazevic2024real_G15], [wang2023task_D8], [huang2024plantography_D14], [chung2022talebrush_W3], [zhang2023visar_W4], [yuan2022wordcraft_W5], [petridis2023anglekindling_W6], [kim2023cells_W9], [benharrak2024writer_W14], [pu2024ideasynth_S1], [liu2024personaflow_S2], [radensky2024scideator_S3], [wang2023popblends_M2], [brade2023promptify_M3], [wang2024lave_M4], [qian2024shape_M5], [chen2024foundation_M8]
Interview-based Methods semi-structured interview inductive & deductive [goodman2022lampost_W1], [reza2024abscribe_W13]
grounded theory [wan2024felt_W8]
unstructured Interview inductive [huang2023causalmapper_G5]
simulated Interview inductive [kocaballi2023conversational_D1]
inductive [blazevic2024real_G15], [brade2023promptify_M3]
observation inductive & deductive [reza2024abscribe_W13]
Observation-based Methods think-aloud inductive [kang2024biospark_G14], [hou2024c2ideas_D13], [chung2022talebrush_W3], [benharrak2024writer_W14], [liu2024personaflow_S2], [wang2024lave_M4]
inductive [kang2024biospark_G14], [kim2024towards_W2], [chung2022talebrush_W3]
screen recordings grounded theory [wan2024felt_W8]
system log inductive [huang2023causalmapper_G5], [pu2024ideasynth_S1], [liu2024personaflow_S2]
Log-based Methods prompts log inductive [shaer2024ai_G1]
chat log grounded theory [ege2024chatgpt_D10]
- [lanzi2023chatgpt_D6]
Survey-based Methods survey inductive [shaer2024ai_G1], [xu2024jamplate_G3], [schuller2024generating_G9], [bernstein2024like_G11], [suh2024luminate_G13]
autoethnography inductive [ghajargar2022redhead_W7]
Other Methods case study method - [lee2024prompt_G18]
@@
Interview-based methods (25 out of 34, or 74%) were the most commonly employed, with semi-structured interviews being the predominant form. Semi-structured interviews allow researchers to explore specific topics of interest while providing flexibility for participants to share their thoughts and experiences. For example, two studies [xu2024jamplate_G3, suh2024luminate_G13] utilized semi-structured interviews to gather user feedback on their proposed LLM-assisted ideation tools. Other forms of interviews, such as unstructured interviews [huang2023causalmapper_G5] and simulated interviews [kocaballi2023conversational_D1], were also used to capture more open-ended and naturalistic user responses.
@@
Observation-based methods (11 out of 34, or 32%) were employed to study user behavior and interactions with LLM-assisted ideation systems in real-time. Techniques such as direct observation [blazevic2024real_G15], think-aloud protocols [kang2024biospark_G14, hou2024c2ideas_D13], and screen recordings [kim2024towards_W2] provided researchers with valuable insights into the usability, user experience, and cognitive processes involved in ideation tasks. These methods allowed researchers to identify patterns, challenges, and opportunities for improvement in the design and implementation of LLM-assisted ideation tools.
@@
Log-based methods (5 out of 34, or 15%) involved the analysis of data generated by the LLM-assisted ideation systems themselves. System logs [huang2023causalmapper_G5, pu2024ideasynth_S1], prompts logs [shaer2024ai_G1], and chat logs [ege2024chatgpt_D10] provided researchers with objective data on user interactions, input patterns, and system responses. By examining these logs, researchers could gain insights into user behavior, preferences, and the effectiveness of different system components in supporting the ideation process.
@@
Survey-based methods (6 out of 34, or 15%), such as questionnaires and surveys, were used to collect user opinions, ratings, and feedback on various aspects of LLM-assisted ideation systems. For example, studies [lanzi2023chatgpt_D6, bernstein2024like_G11] employed surveys to assess user satisfaction, perceived usefulness, and overall experience with the proposed tools and methods.
@@
Other methods (2 out of 34, or 6%), including autoethnography [ghajargar2022redhead_W7] and case study methods [lee2024prompt_G18], were also used to provide rich, contextualized accounts of user experiences and to explore the application of LLM-assisted ideation in specific domains or use cases.
@@
Regarding qualitative data analysis techniques, the majority of the studies (27 out of 34, or 79%) employed inductive approaches, such as thematic analysis or content analysis. Inductive analysis involves identifying themes, patterns, and categories that emerge from the data itself, allowing researchers to develop a deep understanding of user experiences and perspectives. Two studies [goodman2022lampost_W1, reza2024abscribe_W13] combined inductive and deductive analysis, using predefined categories or frameworks to guide the analysis while remaining open to new insights. Additionally, two studies [wan2024felt_W8, ege2024chatgpt_D10] utilized grounded theory, a systematic methodology for developing theories grounded in the data.
@@ 6.3 Quantitative Analysis
4-point Likert scale mean [bhavya2023cam_G16]
- [heyman2024supermind_G2], [xu2024jamplate_G3], [huang2024plantography_D14], [liu2024personaflow_S2]
mean [schuller2024generating_G9], [yuan2022wordcraft_W5]
Likert Scale 5-point Likert scale mean & SD [shaer2024ai_G1], [suh2024luminate_G13], [wang2023task_D8], [hou2024c2ideas_D13], [benharrak2024writer_W14], [banker2024machine_S4], [brade2023promptify_M3]
mean [zhang2023visar_W4]
7-point Likert scale mean & SD [goodman2022lampost_W1], [petridis2023anglekindling_W6], [kim2023cells_W9], [radensky2024scideator_S3], [banker2024machine_S4], [cai2023designaid_M1], [wang2024lave_M4]
absolute rating scale - [chen2024foundation_M8]
NASA-TLX - [reza2024abscribe_W13], [pu2024ideasynth_S1]
7-point Likert scale mean & SD [wang2023popblends_M2], [brade2023promptify_M3]
Consensual Assessment Technique 5-point Likert scale - [chen2024asknaturenet_G10]
IBM Computer Usability Satisfaction Questionnaires 3-point Likert scale - [blazevic2024real_G15]
@@
Likert Scale (20 out of 27, or 74%) was the most commonly used quantitative tool, with variations in the number of points on the scale. Studies employed 4-point [bhavya2023cam_G16], 5-point [shaer2024ai_G1, suh2024luminate_G13], and 7-point [goodman2022lampost_W1, petridis2023anglekindling_W6] Likert scales to measure user perceptions, attitudes, and evaluations of different aspects of the LLM-assisted ideation systems. The choice of scale granularity reflects the researchers’ intended level of precision in capturing user responses.
NASA-TLX (Task Load Index) (5 out of 27, or 19%) was used in five studies [chen2024foundation_M8, reza2024abscribe_W13, pu2024ideasynth_S1, wang2023popblends_M2, brade2023promptify_M3] to assess the subjective workload experienced by users while interacting with the ideation systems. This tool employs both an absolute rating scale and a 7-point Likert scale to measure various dimensions of workload, such as mental demand, physical demand, and effort.
The Consensual Assessment Technique (CAT) (1 out of 27, or 4%) was used in one study [chen2024asknaturenet_G10] to evaluate the creativity of ideas generated with the assistance of an LLM-based system. CAT involves multiple expert judges independently rating the creativity of ideas using a 5-point Likert scale, providing a reliable and validated measure of creativity.
IBM Computer Usability Satisfaction Questionnaires (1 out of 27, or 4%) were employed in one study [blazevic2024real_G15] to assess user satisfaction with an LLM-assisted ideation system. This tool uses a 3-point Likert scale to measure various aspects of usability, such as ease of use, learnability, and overall satisfaction.
@@
Normality Test Shapiro-Wilk Test [shaer2024ai_G1], [kim2023cells_W9]
Independent T-test [kim2023cells_W9], [cai2023designaid_M1]
Mean Comparison Test Paired-sample T-test [yuan2022wordcraft_W5], [pu2024ideasynth_S1], [liu2024personaflow_S2], [radensky2024scideator_S3], [cai2023designaid_M1]
Parametric Test One-tailed T-test [reza2024abscribe_W13]
Welch’s ANOVA [heyman2024supermind_G2]
Analysis of Variance ANOVA [chen2024asknaturenet_G10], [zhang2023visar_W4], [goldi2024intelligent_W10], [wu2023iconshop_M7]
MANOVA [pu2024ideasynth_S1]
Correlation Test Pearson Correlation [shaer2024ai_G1], [liu2024personaflow_S2], [wang2023popblends_M2]
Wilcoxon Rank-Sum Test / Mann-Whitney U Test [schuller2024generating_G9], [goodman2022lampost_W1], [yuan2022wordcraft_W5], [kim2023cells_W9], [goldi2024intelligent_W10]
Rank Test Wilcoxon Signed-Rank Test [pu2024ideasynth_S1], [radensky2024scideator_S3], [banker2024machine_S4], [wang2023popblends_M2], [brade2023promptify_M3]
Non-parametric Test Kruskal-Wallis Test [chen2024asknaturenet_G10]
Multiple Group Comparison Friedman’s Test [hou2024c2ideas_D13]
Post-hoc Test Games-Howell Test [heyman2024supermind_G2]
@@
Novelty Novelty Evaluates whether the idea demonstrates originality and offers a fresh perspective or concept. [chen2024asknaturenet_G10], [bhavya2023cam_G16], [wang2023task_D8], [pu2024ideasynth_S1], [radensky2024scideator_S3], [wu2023iconshop_M7]
Innovativeness Assesses the degree to which the idea introduces innovative approaches that transcend existing norms. [shaer2024ai_G1], [heyman2024supermind_G2]
Uniqueness Examines whether the idea stands out distinctly from other concepts or generated outputs. [wu2023iconshop_M7]
Relevance Relevance Measures the alignment of the idea with the given task, topic, or research objectives. [shaer2024ai_G1], [pu2024ideasynth_S1], [banker2024machine_S4]
Diversity Variety Examines the breadth of perspectives or options presented by the idea, emphasizing its diversity. [chen2024asknaturenet_G10], [chung2022talebrush_W3]
Functionality and Usability Usability Examines whether the idea is intuitive and practical, facilitating ease of adoption and application. [wang2023task_D8], [chung2022talebrush_W3]
Feasibility Determines whether the idea is realistically implementable in practical scenarios. [wang2023task_D8], [pu2024ideasynth_S1]
Efficiency Assesses the idea’s potential to achieve objectives effectively within resource constraints. [chung2022talebrush_W3]
Functionality Evaluates the idea’s capacity to address specific functional needs or solve targeted problems. [wang2023task_D8]
Impact and Inspiration Impact Measures the potential of the idea to effect significant change or address major challenges. [pu2024ideasynth_S1], [banker2024machine_S4]
Insightfulness Assesses whether the idea provides deep, unique insights that enhance understanding of a problem. [shaer2024ai_G1]
Inspiration Evaluates the idea’s ability to stimulate further creative thinking or ideation. [chung2022talebrush_W3]
Meaningfulness Considers whether the idea is contextually significant and capable of evoking value or resonance. [bhavya2023cam_G16]
Logical Coherence Plausibility Determines whether the idea is logical, realistic, and consistent with common sense or knowledge. [banker2024machine_S4]
Clarity and Specificity Clarity Assesses whether the idea is clearly articulated and unambiguous, enhancing its comprehensibility. [banker2024machine_S4]
Specificity Evaluates the level of detail and precision in the idea, ensuring actionable outcomes. [pu2024ideasynth_S1]

### v3

@@ 6.2 Qualitative Analysis
inductive [83], [87], [61], [90], [78], [62], [50], [54], [47], [49], [44], [46], [67], [68], [69], [58], [59], [66], [64], [60]
Interview-based Methods semi-structured interview inductive & deductive [56], [45]
grounded theory [48]
unstructured Interview inductive [74]
simulated Interview inductive [35]
inductive [90], [59]
observation inductive & deductive [45]
Observation-based Methods think-aloud inductive [61], [40], [50], [46], [68], [66]
inductive [61], [43], [50]
screen recordings grounded theory [48]
system log inductive [74], [67], [68]
Log-based Methods prompts log inductive [95]
chat log grounded theory [37]
- [80]
Survey-based Methods survey inductive [95], [83], [89], [77], [87]
autoethnography inductive [51]
Other Methods case study method - [76]
@@
Interview-based methods (25 out of 34, or 74%) were the most commonly employed, with semi-structured interviews being the predominant form. Semi-structured interviews allow researchers to explore specific topics of interest while providing flexibility for participants to share their thoughts and experiences. For example, two studies [83, 87] utilized semi-structured interviews to gather user feedback on their proposed LLM-assisted ideation tools. Other forms of interviews, such as unstructured interviews [74] and simulated interviews [35], were also used to capture more open-ended and naturalistic user responses.
@@
Observation-based methods (11 out of 34, or 32%) were employed to study user behavior and interactions with LLM-assisted ideation systems in real-time. Techniques such as direct observation [90], think-aloud protocols [61, 40], and screen recordings [43] provided researchers with valuable insights into the usability, user experience, and cognitive processes involved in ideation tasks. These methods allowed researchers to identify patterns, challenges, and opportunities for improvement in the design and implementation of LLM-assisted ideation tools.
@@
Log-based methods (5 out of 34, or 15%) involved the analysis of data generated by the LLM-assisted ideation systems themselves. System logs [74, 67], prompts logs [95], and chat logs [37] provided researchers with objective data on user interactions, input patterns, and system responses. By examining these logs, researchers could gain insights into user behavior, preferences, and the effectiveness of different system components in supporting the ideation process.
@@
Survey-based methods (6 out of 34, or 15%), such as questionnaires and surveys, were used to collect user opinions, ratings, and feedback on various aspects of LLM-assisted ideation systems. For example, studies [80, 77] employed surveys to assess user satisfaction, perceived usefulness, and overall experience with the proposed tools and methods.
@@
Other methods (2 out of 34, or 6%), including autoethnography [51] and case study methods [76], were also used to provide rich, contextualized accounts of user experiences and to explore the application of LLM-assisted ideation in specific domains or use cases.
@@
Regarding qualitative data analysis techniques, the majority of the studies (27 out of 34, or 79%) employed inductive approaches, such as thematic analysis or content analysis. Inductive analysis involves identifying themes, patterns, and categories that emerge from the data itself, allowing researchers to develop a deep understanding of user experiences and perspectives. Two studies [56, 45] combined inductive and deductive analysis, using predefined categories or frameworks to guide the analysis while remaining open to new insights. Additionally, two studies [48, 37] utilized grounded theory, a systematic methodology for developing theories grounded in the data.
@@ 6.3 Quantitative Analysis
4-point Likert scale mean [85]
- [84], [83], [62], [68]
mean [89], [47]
Likert Scale 5-point Likert scale mean & SD [95], [87], [78], [40], [46], [70], [59]
mean [54]
7-point Likert scale mean & SD [56], [49], [44], [69], [70], [57], [66]
absolute rating scale - [60]
NASA-TLX - [45], [67]
7-point Likert scale mean & SD [58], [59]
Consensual Assessment Technique 5-point Likert scale - [86]
IBM Computer Usability Satisfaction Questionnaires 3-point Likert scale - [90]
@@
Likert Scale (20 out of 27, or 74%) was the most commonly used quantitative tool, with variations in the number of points on the scale. Studies employed 4-point [85], 5-point [95, 87], and 7-point [56, 49] Likert scales to measure user perceptions, attitudes, and evaluations of different aspects of the LLM-assisted ideation systems. The choice of scale granularity reflects the researchers’ intended level of precision in capturing user responses.
NASA-TLX (Task Load Index) (5 out of 27, or 19%) was used in five studies [60, 45, 67, 58, 59] to assess the subjective workload experienced by users while interacting with the ideation systems. This tool employs both an absolute rating scale and a 7-point Likert scale to measure various dimensions of workload, such as mental demand, physical demand, and effort.
The Consensual Assessment Technique (CAT) (1 out of 27, or 4%) was used in one study [86] to evaluate the creativity of ideas generated with the assistance of an LLM-based system. CAT involves multiple expert judges independently rating the creativity of ideas using a 5-point Likert scale, providing a reliable and validated measure of creativity.
IBM Computer Usability Satisfaction Questionnaires (1 out of 27, or 4%) were employed in one study [90] to assess user satisfaction with an LLM-assisted ideation system. This tool uses a 3-point Likert scale to measure various aspects of usability, such as ease of use, learnability, and overall satisfaction.
@@
Normality Test Shapiro-Wilk Test [95], [44]
Independent T-test [44], [57]
Mean Comparison Test Paired-sample T-test [47], [67], [68], [69], [57]
Parametric Test One-tailed T-test [45]
Welch’s ANOVA [84]
Analysis of Variance ANOVA [86], [54], [55], [65]
MANOVA [67]
Correlation Test Pearson Correlation [95], [68], [58]
Wilcoxon Rank-Sum Test / Mann-Whitney U Test [89], [56], [47], [44], [55]
Rank Test Wilcoxon Signed-Rank Test [67], [69], [70], [58], [59]
Non-parametric Test Kruskal-Wallis Test [86]
Multiple Group Comparison Friedman’s Test [40]
Post-hoc Test Games-Howell Test [84]
@@
Novelty Novelty Evaluates whether the idea demonstrates originality and offers a fresh perspective or concept. [86], [85], [78], [67], [69], [65]
Innovativeness Assesses the degree to which the idea introduces innovative approaches that transcend existing norms. [95], [84]
Uniqueness Examines whether the idea stands out distinctly from other concepts or generated outputs. [65]
Relevance Relevance Measures the alignment of the idea with the given task, topic, or research objectives. [95], [67], [70]
Diversity Variety Examines the breadth of perspectives or options presented by the idea, emphasizing its diversity. [86], [50]
Functionality and Usability Usability Examines whether the idea is intuitive and practical, facilitating ease of adoption and application. [78], [50]
Feasibility Determines whether the idea is realistically implementable in practical scenarios. [78], [67]
Efficiency Assesses the idea’s potential to achieve objectives effectively within resource constraints. [50]
Functionality Evaluates the idea’s capacity to address specific functional needs or solve targeted problems. [78]
Impact and Inspiration Impact Measures the potential of the idea to effect significant change or address major challenges. [67], [70]
Insightfulness Assesses whether the idea provides deep, unique insights that enhance understanding of a problem. [95]
Inspiration Evaluates the idea’s ability to stimulate further creative thinking or ideation. [50]
Meaningfulness Considers whether the idea is contextually significant and capable of evoking value or resonance. [85]
Logical Coherence Plausibility Determines whether the idea is logical, realistic, and consistent with common sense or knowledge. [70]
Clarity and Specificity Clarity Assesses whether the idea is clearly articulated and unambiguous, enhancing its comprehensibility. [70]
Specificity Evaluates the level of detail and precision in the idea, ensuring actionable outcomes. [67]